Teall Development Company hired you as a consultant to help them estimate its cost of capital. You have been provided with the following data: D1 = $1.45; P0 = $26.00; and g = 6.50% (constant). Based on the DCF approach, what is the cost of equity from retained earnings?

Group of answer choices

12.68%

11.84%

12.08%

10.63%

9.78%
Answer #1

Ans 12.08%

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
26 = 1.45 / (Ke - 6.50%)
Ke - 6.50% = 1.45 / 26
Ke - 6.50% = 5.58%
Ke = 12.08%